OpenStudy (anonymous):

Help please, If tan(x)=2 and tan(x+y)=135 Find tan(y)

jhonyy9 (jhonyy9):

sin(x+y) tan(x+y) = ---------- = cos(x+y) hope you know these formulas and will can continue it

OpenStudy (anonymous):

Ok , I'll try that, so far I've only tried the tangent addition identity which is tan(x)+tan(y) ------------ 1-tan(x)tan(y)

OpenStudy (anonymous):

Do you think that would work if I took that route?

OpenStudy (mathmale):

Since the only other info that you were given is "tan x =2," yes, I think you should stay with the sum formula for the tangent.

jimthompson5910 (jim_thompson5910):

\[\Large \tan(x+y) = \frac{\tan(x)+\tan(y)}{1-\tan(x)\tan(y)}\] \[\Large 135 = \frac{2+\tan(y)}{1-2\tan(y)}\] \[\Large 135 = \frac{2+z}{1-2z} \ ... \ \text{Let } z = \tan(y)\] now solve for z

OpenStudy (anonymous):

All right thanks guys :)